TEMPLATE FOR REU PAPERS

YOUR NAME

Abstract. Give a brief description of your paper here

Contents

1. What does the table of contents command do? The table of contents command will automatically make a contents. You must

run tex at least twice for this to work.

2. How do the environment commands work?

The command name is indicated as in “defn” above. To get a definition, you type: Definition 2.1. This is how to define a definition.

And for a theorem and its proof you would type: Theorem 2.2. This is the statement of a theorem. Proof. And this shows that the statement is correct. □

Note that the numbering is taken care of automatically. Equations get numbered in sequence with statements, as for example

(2.3)  e = mc^{2}

I’m hoping you have xypic available, and so have included a relevant usepackage. It works like this.

\xymatrix{F X \ar[r]^{F f} \ar[d]_{η_{X}} & F Y \ar[d]^{η_{Y}} \\ GX \ar[r]_{Gf} & GY}

References

[1] J. P. May. A Concise Course in Algebraic Topology. University of Chicago Press. 1999.

Date: DEADLINE AUGUST 22, 2008.